### v1

# HGCN(O): A Self-Tuning GCN HyperModel Toolkit for Outcome Prediction in Event-Sequence Data

Abstract
We propose HGCN(O), a self-tuning toolkit using Graph Convolutional Network (GCN) models for event sequence prediction. Featuring four GCN architectures (O-GCN, T-GCN, TP-GCN, TE-GCN) across the GCNConv and GraphConv layers, our toolkit integrates multiple graph representations of event sequences with different choices of node- and graph-level attributes and in temporal dependencies via edge weights, optimising prediction accuracy and stability for balanced and unbalanced datasets. Extensive experiments show that GCNConv models excel on unbalanced data, while all models perform consistently on balanced data. Experiments also confirm the superior performance of HGCN(O) over traditional approaches. Applications include Predictive Business Process Monitoring (PBPM), which predicts future events or states of a business process based on event logs.
keywords:
Business Process Monitoring , Graph Convolutional Networks , Outcome-Oriented Event-Sequence Prediction , Hyper Model
journal: Knowledge-Based Systems
[…]
[L1]organization=College of Computing and Mathematical Sciences, Khalifa University, addressline= P.O. Box 127788, city=Abu Dhabi, country=United Arab Emirates
[…]
[L2]organization=Center for Cyber-Physical Systems, Khalifa University, addressline= P.O. Box 127788, city=Abu Dhabi, country=United Arab Emirates
[…]
[L3]organization=Department of Computer Science, University of Milan, addressline= Via Festa del Perdono 7, postcode =20122, city=Milano, state=MI, country=Italy
[…]
[L4]organization=Secure Service-oriented Architectures Research Lab, University of Milan, addressline= Via Celoria 18, postcode =20133, city=Milano, state=MI, country=Italy
Highlights
  • 1.
    Self-tuning GCN toolkit for outcome prediction in Predictive Business Processes
  • 2.
    Edge-weighted GCNs capture intricate temporal and structural process dependencies
  • 3.
    Multiple GCN variants address balanced and imbalanced datasets effectively
  • 4.
    Benchmark for future PBPM research with diverse graph-based input representations
1 Introduction
Event sequence prediction consists of predicting future events and outcomes based on sequences of past events. Event prediction using Machine Learning (ML) is a well-developed research area [1]. A significant application domain is Predictive Business Process Monitoring (PBPM) [2], where the goal is to predict the outcome of a business process instance from incomplete executions [3, 4, 5]. Outcome prediction is usually framed as a classification problem, where each instance of the process is classified into one of several possible outcomes. For example, predicting whether a customer order will be delivered on time can be stated as a binary classification problem where the classifier learns to assign one of the predefined classes (on time vs. delayed) to each process instance based on the sequence of events that lead to the prediction point [6]. Basic ML techniques have been successfully applied to PBPM [7], notably Random Forest (RF) [8], XGBoost [9], and Decision Trees [10, 11, 12]. Compared to support vector machines (SVM), RF and boosted trees demonstrated superior performance in the prediction of the outcomes of business process consisting of periodic or near-periodic events [13].
The capability of deep-ML models to learn complex patterns in sequential data suggested using them to predict the outcomes of processes based on intricate, non-periodic event sequences. Some pioneering works [14] applied Recurrent Neural Networks to predict process outcomes, paving the way to the adoption of long short-term memory (LSTM) and other stateful deep neural network classifiers [15]. Based on this foundation, Wang et al. [16] employed attention-based LSTMs to capture complex temporal dependencies, outperforming traditional methods in benchmark evaluations. In addition, convolutional neural network approaches have emerged, utilizing image representations of process traces to work in a latent space highly suitable for prediction of multi-class labels [6].
In recent years, graph embeddings have become popular as a method to capture aspects of sequential order and provide them to ML algorithms downstream [17]. Graph-based ML models offer many advantages in outcome prediction, particularly in capturing many to many relationships between events within a process trace. Graph Convolutional Networks (GCNs) support a multilevel feature structure, which can be used to integrate information on individual events (features at the node level) and information on the overall process (features at the graph level), making them well suited for representing intricate PBPM data. Incorporating edge weights allows for modeling complex temporal relations and dependencies between events. However, despite their success in other domains, graph-based ML models remain relatively unexplored in PBPM [2]. Business processes often correspond to complex dynamic graph structures that make manual hyperparameter selection and GCN model tuning prohibitive [18].
To address this gap, we propose HGCN(O), a toolkit based on self-tuning hypermodels with diverse input structures, optimized for balanced and unbalanced data sets. Our framework adapts seamlessly to various data characteristics and prediction requirements within PBPM.
This paper presents what we believe to be the first application of hyper-GCN models to outcome prediction tasks in PBPM, with several key contributions. First, we describe a novel toolkit that supports the entire life cycle of ML models from attribute encoding to model implementation, providing a benchmark for future graph-based PBPM research. Second, we introduce a self-tuning mechanism that dynamically optimizes hyperparameters for each GCN model, ensuring adaptability to both balanced and unbalanced datasets. Third, we propose a flexible input structure that includes multiple variants of the GCN model (O-GCN, T-GCN, TP-GCN, and TE-GCN), each tailored to represent different types and levels of encoded and embedded attributes at both the node (event) and graph (process) levels. Our models use edge weights to represent elapsed time between activities, allowing models to differentiate between single-event durations and inter-event timing, decreasing training time, and improving prediction accuracy. Validated on benchmark datasets, our toolkit demonstrates superior performance, establishing a robust, scalable framework for real-time outcome prediction in PBPM applications.
The paper is structured as follows. Section 2 presents related work on graph representation. Section 3 introduces the basic mechanisms of GCNs. Section 4 outlines the graph representation of the PBPM data. Section 5 details the architectures of our GCN hyper models fitted to specific scenarios. Section 6 presents the experimental results for multiple datasets. Finally, Section 7 summarizes the findings and outlines our future research directions.
2 Related Work
Graphs are a popular language for representing tabular data of all kinds. Graph representation makes choices in encoding data values and relationships into graph parameters such as nodes and edge attributes, while graph representation learning consists of processing graph representations and transforming them into a form suitable for training ML models. The goal of graph representation learning is to define vector representations of graph entities (e.g. graph nodes, edges, and sub-graphs) to facilitate node classification, link prediction, community detection, and others. Graph representation learning plays an important role as it could significantly improve the performance of downstream ML tasks. Over the decades, several techniques have been proposed to compute fixed-length vectors encoding graphs, including graph kernels, matrix factorization models, shallow models, deep neural network models, and non-Euclidean models [19, 20, 21, 22]. Graph kernels are kernel functions that measure the similarity between graphs [23] and between the nodes of a given graph. However, graph kernels do not scale well, as computing graph kernel functions is an NP-hard class [24]. Other approaches rely on matrix factorisation methods to decompose the graph adjacency matrix into products of smaller matrices and then learn vector embeddings that fit each of them. Proximity matrix factorisation models have successfully handled large graphs [25, 26] but suffer from scalability problems on huge datasets due to the computational complexity of factorising large dense matrices. The second half of the last decade saw the emergence of shallow graph embedding models such as DeepWalk [27], Node2Vec [28], and LINE [29]. These methods rely on techniques inspired by natural language processing, such as the skip-gram model, to learn vector embeddings of graph nodes by sampling node neighborhoods through random walks. Unlike deep models, they do not use multi-layer neural networks, but optimise the embeddings directly. These methods are scalable to large graphs and allow the representation of graph entities in low-dimensional vector spaces. They differ mainly in how they sample neighborhoods and preserve structural properties of the graph, such as proximity or community structure. However, these models struggle to generalise to unseen nodes or graphs, and often fail to effectively incorporate rich node features or dynamic graph changes [30].
The advent of deep learning led to a new research perspective, introducing graph neural networks as a powerful framework for learning graph embeddings. Models such as recurrent GNNs (R-GNNs) and their successors have demonstrated significant expressiveness, overcoming some limitations of shallow models, such as their inability to exploit rich node features or generalise to unseen graphs [31, 32, 33]. Most R-GNN models learn node embeddings by sharing weights across their recurrent layers, a design choice that simplifies training but can limit the model’s ability to distinguish between local and global graph structures. To address such limitations, graph autoencoder models have been proposed that exploit the principles of the original autoencoder architecture to encode graph representations in a way that can capture both structural and feature-level information [34, 35]. Graph auto-encoders consist of two main layers: encoder layers, which take the adjacency matrix as input and reduce its dimensionality to generate node embeddings, and decoder layers, which reconstruct the adjacency matrix from these embeddings. Inspired by the transformer architecture widely used in natural language processing, graph transformer models adapt this architecture for graph-based tasks [36, 37]. Early graph transformer models focused primarily on learning tree-structured or hierarchical graphs [38, 39]. Modern graph transformer models extend their capabilities by encoding node positions using both relative and absolute position encodings. This advance allows them to effectively learn complex graph structures, although they perform best on graphs with some degree of structural regularity [40].
In this paper, we introduce a novel approach using convolutional operators with different weights in each hidden layer, a technique that shows promise in capturing and distinguishing local and global graph structures. This concept has served as a catalyst for extensive research on GCNs, driving advances in their design, implementation, and application in diverse domains, including social network analysis [29], molecular biology [41], and neuroanatomy [42]. Specifically, we apply this idea to outcome prediction in PBPM, where GCNs have been shown to improve prediction accuracy [43, 44, 45, 46]. However, the existing literature primarily explores a narrow range of strategies for encoding event sequences, often focusing on predefined feature extraction methods or simplistic sequence representations that fail to capture the complex temporal dependencies and contextual relationships inherent in many real-world business processes [14, 47, 6, 48]. Our work addresses this gap by introducing a self-tuning mechanism that dynamically adapts to the specific characteristics of the business process under analysis. A detailed introduction to GCNs is given in the next section.
3 Graph Convolutional Networks
GCNs learn from graph-structured data by aggregating information from neighbouring nodes through convolutional operations. In GCNs, each layer updates its node representations based on the graph structure and node characteristics. In this paper, we rely on two types of graph convolutional layers: GCNConv and GraphConv, which differ in their mathematical formulation and in their treatment of graph data. GCNConv, introduced by Kipf and Welling [49], uses a renormalised adjacency matrix to normalise the aggregation of neighbour information [49]. This representation is formulated and adapted to our study as follows.
$\mathbb{V}^{(l+1)}=\sigma\left(\tilde{D}^{-\frac{1}{2}}\tilde{A}_{w}\tilde{D}^{-\frac{1}{2}}\mathcal{G}^{(l)}W^{(l)}\right),$ (1)
where $A_{w}$ is the original weighted adjacency matrix, where each entry $A_{w_{(i\rightarrow i+1)}}$ represents the edge weight $w_{(i\rightarrow i+1)}$ between node $i$ and node $i+1$. $\tilde{D}$ is the degree matrix calculated from $\tilde{A}_{w}$. $\tilde{A}_{w}=A_{w}+I$ is a weighted adjacency matrix with added self-loops, allowing nodes to aggregate information from themselves and from their neighbors. $\mathcal{G}^{(l)}=(\mathbb{V}^{(l)},\mathbb{E},\mathbb{W})$ represents the graph data at layer $l$, including the characteristics of the nodes, the edge indices, and the edge weights, respectively. Symmetrical normalization $\tilde{D}^{-\frac{1}{2}}\tilde{A}_{w}\tilde{D}^{-\frac{1}{2}}$ is crucial for handling varying degrees of nodes and weighted relationships. $W^{(l)}$ is the learnable weight matrix and $\sigma$ is the activation function. This normalization stabilizes feature scaling during message passing, making GCNConv very effective, particularly in semi-supervised tasks where node degree variance can impact learning. In contrast, GraphConv [50] uses the following adapted operation:
$\mathbb{V}^{(l+1)}=\sigma\left(D_{w}^{-1}A_{w}\mathcal{G}^{(l)}W^{(l)}\right)$ (2)
where $D$ is the degree matrix computed from $A_{w}$. Unlike GCNConv, GraphConv does not apply symmetric normalization. Instead, the weighted degree matrix $D_{w}^{-1}$ scales the aggregation of characteristics directly. This formulation enables the model to handle edge weights, ensuring that neighbor contributions are appropriately weighted during the convolution process. Although GraphConv offers greater flexibility, it can sacrifice stability when handling graphs with significant degree variance. Its simpler formulation can however be advantageous for uniform graph structures, though the absence of adjacency normalization makes neighbor information aggregation more critical.
The initial formulations of GCNConv and GraphConv [49, 50] did not support edge weights, which are crucial in many real-world applications. In this study, we use the PyTorch Geometric GCNConv implementation [51, 52], which extends graph convolutional models to enable weighted message passing, capturing the relationships between nodes. Given the importance of edge weights in our analysis, we do not consider GCN variants such as SageConv [53], ClusterGCNConv [54] and EdgeConv [55] that do not support edge attributes and rely solely on the adjacency matrix for feature aggregation. Since edge weights represent relational importance between nodes, their omission limits the applicability of these models to our work.
Our toolkit relies on the two foundational GCN models, GCNConv and GraphConv that provide a solid basis for processing dynamic graph structures. Although we considered alternative approaches such as Graph Attention Networks (GAT) [56] and Graph Isomorphism Networks (GIN) [57], we ultimately excluded them from our study due to some specific characteristics that do not align with our objectives. Indeed, GAT’s over-reliance on attention mechanisms increases its computational overhead, making the model prone to scalability problems when dealing with huge graphs, which are common in PBPM. GIN’s high sensitivity to even small variations in the graph structure is hardly necessary for PBPM purposes. By focusing on GCNConv and GraphConv, we achieve an optimal balance of performance and scalability.
4 Hierarchical Graph Attribute Representation and Encoding
4.1 Node and Graph Attributes Notation and Encoding
The analysis of graph representations of sequential data occurs at two hierarchical levels: the node level and the graph level. Let $X_{i}$ denote a node representing an individual event, while $G_{j}$ encapsulates a complete sequence of events. The notation $X_{i}\in G_{j}$ indicates that the event $X_{i}$ is part of the sequence $G_{j}$. Attributes ($F^{a}_{j},F^{b}_{j},F^{c}_{j},$ etc.) at the graph level capture characteristics of the entire sequence $G_{j}$, providing a comprehensive view of the individual graph.
[…]
As far as typing is concerned, attributes are categorized into two main types: categorical attributes, encoded using one-hot encoding, and numerical attributes, processed through min-max scaling. When dealing with irrelevant values associated with specific attributes $B_{i}$ that apply only to certain nodes, different strategies are used depending on the attribute type. For numerical attributes, these values are replaced with the median to mitigate the impact of skewness in the data. For categorical attributes, irrelevant values are encoded as 1-1- 1, which also serves as padding. During the training process, padding values are masked to ensure that they do not affect the performance of the model.
For each graph node $N_{i}$, we concatenate all encoded attributes into a unified composite vector, denoted as: $\mathbf{v}_{N_{i}}=\left[\mathcal{A}_{i},B_{i},U_{i}\right]$. For each sequence graph $G_{j}$, we represent the sequence attributes as: $\mathbf{v}_{G_{j}}=\left[F_{j}\right]$.
4.2 Edge Weight Representation
We define the edge weights for the graph as a one-dimensional vector representing the time differences (in seconds) between the start times of “activities” in consecutive nodes, calculated as $w_{(i\rightarrow i+1)}=T^{s}_{i+1}-T^{s}_{i}$. If “activities” share the same start time, the edge weight is set to 0, capturing the simultaneous nature of these nodes. This approach enables the model to readily capture temporal dependencies while also distinguishing between simultaneous and sequential relationships among nodes, thereby improving predictive accuracy. To ensure consistency during training, we apply min-max scaling to normalize these weights to a range between 0 and 1.
4.3 Graph Representation Construction
Given an event sequence, we construct a graph representation where each node is represented by a preprocessed vector $\mathbf{v}_{N_{i}}\in\mathbb{R}^{d_{N}}$, with $d_{N}$ denoting the dimensionality of the node vector. The number of nodes $n$ is determined by counting the valid entries in each graph. The matrix of node vectors for graph $G_{j}$ is defined as $\mathbb{V}_{N(G_{j})}\in\mathbb{R}^{n\times d_{N}}$, where $\mathbb{V}_{N(G_{j})}=[\mathbf{v}_{N_{1}},\ldots,\mathbf{v}_{N_{n}}]^{\intercal}$. The edge index tensor $\mathbb{E}_{G_{j}}\in\mathbb{R}^{2\times(n-1)}$ is generated to connect consecutive nodes, with each edge $e_{N_{(i\rightarrow i+1)}}$ defined as $e_{N_{(i\rightarrow i+1)}}=\mathbb{E}_{G_{j}}^{i}$ for $i=1,\ldots,n-1$. The corresponding edge weights are stored in the tensor $\mathbb{W}_{G_{j}}\in\mathbb{R}^{n-1}$, where $w_{(i\rightarrow i+1)}=\mathbb{W}_{G_{j}}^{i}$ for $i=1,\ldots,n-1$. Finally, we create a graph data object $\mathcal{G}_{G_{j}}=(\mathbb{V}_{N(G_{j})},\mathbb{E}_{G_{j}},\mathbb{W}_{G_{j}})$, that includes node attributes, edge indices, and edge weights.
Some studies [58] define the edge direction between graph nodes based on the completion time of the sequence events. In contrast, our approach uses the start time of the activities to define the edge weights. To handle scenarios where an activity’s duration may extend past the start of the following one, we include a duration attribute (as previously defined) within each node representation. This design enables our model to capture nuanced temporal relationships between events, providing a more accurate analysis of sequential data.
5 GCN HyperModels Architectures and Optimization
This section proposes four self-tuning hyper-model architectures based on two types of GCN models: GraphConv and GCNConv. Each architecture is specifically designed for a unique input pipeline and configuration, optimizing performance across diverse datasets.
5.1 One-Level GCN (O-GCN)
For OnelevelGCN (O-GCN) model, we integrate node-level and graph-level vectors to create a unified input representation for each node. The graph-level vector $\mathbf{v}_{G_{j}}\in\mathbb{R}^{d_{G_{j}}}$ is expanded to match the number of nodes, yielding $\mathbf{v}_{G_{j}}^{\prime}\in\mathbb{R}^{n\times d_{G_{j}}}$, where $n$ is the number of nodes. This expansion is achieved by repeating $\mathbf{v}_{G_{j}}$ across all nodes. The final input representation is formed by concatenating the node vectors $\mathbb{V}_{N_{(}G_{j})}$ with the expanded graph-level vector, resulting in $\mathbb{V}_{N(G_{j})}^{\prime}=[\mathbb{V}_{N(G_{j})},\mathbf{v}_{G_{j}}^{\prime}]\in\mathbb{R}^{n\times(d_{N}+d_{G_{j}})}$. Each node is then represented by the combined vector $\mathbf{v}_{N_{i}}^{\prime}=\mathbb{V}_{N(G_{j})}^{\prime i}$ for $i=1,\dots,n$. The graph data object is updated to $\mathcal{G}_{G_{j}}^{\prime}=(\mathbb{V}_{N(G_{j})}^{\prime},\mathbb{E}_{G_{j}},\mathbb{W}_{G_{j}})$, which serves as input to the model. This approach effectively integrates local node attributes with global graph-level information, making it particularly suitable for GCN layers.
The graph object $\mathcal{G}_{G_{j}}^{\prime}$ is initially processed through a series of GCNConv or GraphConv layers, yielding $\mathcal{G}_{G_{j}}^{\prime L}=(\mathbb{V}_{N(G_{j})}^{\prime L},\mathbb{E}_{G_{j}},\mathbb{W}_{G_{j}})$, where $\mathbb{V}_{N(G_{j})}^{\prime L}$ represents the node features at the final layer $L$. A pooling operation aggregates node representations $\mathbb{V}_{N(G_{j})}^{\prime L}$ into a graph-level embedding $\mathbf{z}_{G_{j}}$. This embedding is passed through a stack of fully connected layers, producing $\mathbf{z}_{G_{j}}^{(fc)}$ at final layer $fc$. The classification output $\hat{y}_{G_{j}}$ is classically obtained by applying a linear classifier followed by a soft-max activation.
5.2 Two-Level GCN (T-GCN)
The Two-Level GCN (T-GCN) model initiates by passing the graph object $\mathcal{G}_{G_{j}}$ (containing only node-level features) through a series of GCNConv or GraphConv layers. This results in an updated graph object $\mathcal{G}_{G_{j}}^{L}=(\mathbb{V}_{N(G_{j})}^{L},\mathbb{E}_{G_{j}},\mathbb{W}_{G_{j}})$. A pooling operation is then applied to aggregate the node representations into a graph-level embedding, $\mathbf{z}_{G_{j}}$. Simultaneously, preprocessed graph-level attribute vectors, $\mathbf{v}_{G_{j}}$ are passed through a series of dense layers, producing $\mathbf{v}_{G_{j}}^{d}$ at final layer $d$. The combined representation is then constructed as: $\mathcal{Z}_{G_{j}}=\textit{Concatenate}(\mathbf{z}_{G_{j}},\mathbf{v}_{G_{j}}^{d})$.
Similar to the O-GCN, $\mathcal{Z}_{G_{j}}$ is passed through a series of fully connected layers, culminating in a final output layer to generate the classification output $\hat{y}_{G_{j}}$.
5.3 Two-Level Pseudo-Embedding GCN (TP-GCN)
The Two-Level Pseudo-Embedding GCN (TP-GCN) model incorporates an additional input in the form of a pseudo-embedding matrix. This matrix consists of preprocessed embeddings derived from specific node attributes, providing an alternative feature space to complement the raw node features. Techniques such as node2vec and DeepWalk [27] [28] are commonly used to generate a fixed-dimensional embedding matrix, where each row represents the learned representation of a node based on its structural connectivity within the graph. In this study, we employ a pseudo-embedding duration matrix [59] based ion duration binning, utilizing a Term Frequency Inverse Document Frequency (TF-IDF) approach to capture relationships between nodes and their associated duration attributes (see Algorithm 1). TF-IDF is a popular text retrieval technique for computing how relevant a term is to a specific document belonging to a corpus. The relevance increases proportionally to the number of times the term appears in the document but is compensated for by the frequency of the term in the corpus.
Algorithm 1 Pseudo-Embedding Duration Bin Matrix
0: Set of activity (node) ${\mathcal{A}_{i}}$; duration value $T^{d}_{i}$ for each $\mathcal{A}_{i}$; set of graphs ${G_{j}}$.
0: Pseudo-embedding matrices with each $\mathcal{A}_{i}$ represented as a vector $\mathbf{v}_{bin_{i}}$ for all $G_{j}$.
1: Initialize cut-off value $T_{\text{cut}}$ and number of quantile bins $\mathcal{N}_{\mathring{b}}$.
2:repeat
3:  for each event $\mathcal{A}i$ do
4:   if $T^{d}_{i}<T{\text{cut}}$ then
5:    Assign $T^{d}_{i}$ to unique bin $b$.
6:   else
7:    Calculate quantile bins $\mathring{b}$ based on $\mathcal{N}_{\mathring{b}}$.
8:    Remove duplicates and update $\mathring{b}$ for full range coverage.
9:    Assign bins $T^{d_{\mathring{b}}}i$.
10:   end if
11:   Assign duration bin $T^{d_{b^{\prime}}}_{i}$ to $\mathcal{A}_{i}$.
12:  end for
13:  Calculate bin frequencies $\{f_{b}^{\prime}\}$, where $\{f_{b}^{\prime}\}=\{f_{b},f_{\mathring{b}}\}$.
14:  if any $f_{\mathring{b}}\approx$ any $f_{\mathring{b}}$ then
15:   BREAK
16:  else
17:   Update $T_{\text{cut}}$ and $\mathcal{N}_{\mathring{b}}$.
18:  end if
19:until stopping condition is met
20: Extract unique combinations $(\mathcal{A}i,T^{d_{b^{\prime}}})$.
21: Treat each $(\mathcal{A}i,T^{d_{b^{\prime}}})$ as a term.
22: Construct corpus from ${(\mathcal{A}i,T^{d_{b^{\prime}}})}$.
23:for each graph $G_{j}$ do
24:  Treat graph $G_{j}$ as a document.
25:  Construct tf-idf matrix for $G_{j}$:
26:  for each $(\mathcal{A}_{i},T^{d_{b^{\prime}}})$ do
[…]
28:  end for
29:  Construct tf-idf matrix with columns for $T^{d_{b^{\prime}}}$ and rows for $\mathcal{A}_{i}$.
30:end for
31:return  Pseudo-embedding matrices with each $\mathcal{A}_{i}$ as vector $\mathbf{v}_{bin_{i}}$ for all $G_{j}$.
Our duration matrix, denoted as $\mathbb{V}_{\mathring{N}(G_{j})}$, is built by applying the same concept to the graph nodes. It contains node vectors $\mathbf{v}_{\mathring{N}_{i}}$, where $\mathbf{v}_{\mathring{N}_{i}}=\mathbb{V}_{\mathring{N}(G_{j})}^{i}$ for $i=1,\dots,n$. Consequently, an additional graph object $\mathcal{G}_{\mathring{G}_{j}}$ is created, sharing the same edge indices and weights as $\mathcal{G}_{G_{j}}$. It is expressed as $\mathcal{G}_{\mathring{G}_{j}}=(\mathbb{V}_{\mathring{N}(G_{j})},\mathbb{E}_{G_{j}},\mathbb{W}_{G_{j}})$.
Both $\mathcal{G}_{G_{j}}$ and $\mathcal{G}_{\mathring{G}_{j}}$ are processed through separate stacks of GCN layers, producing outputs $\mathcal{G}_{G_{j}}^{L}$ and $\mathcal{G}_{\mathring{G}_{j}}^{L}$, respectively. These outputs are then concatenated to form a unified feature vector, $\zeta_{G_{j}}$, where $\zeta_{G_{j}}=\textit{Concatenate}(\mathbb{V}_{N({G}j})^{L},\mathbb{V}_{\mathring{N}({G}_{j}})^{L})$. The graph object is updated as $\mathcal{G}_{\ddot{G}_{j}}=(\zeta_{G_{j}},\mathbb{E}_{G_{j}},\mathbb{W}_{G_{j}})$. The updated graph is processed through a series of GCN layers, followed by a pooling operation to generate a graph-level embedding $\mathbf{z}_{G_{j}}$. Similar to T-GCN, the subsequent steps involve incorporating the graph-level vector $\mathbf{v}_{G_{j}}$ and its dense layers, concatenating vectors, passing the combined representation through fully connected layers and applying a final linear classifier.
5.4 Two-Level Embedding GCN (TE-GCN)
[…]
G
j
[…]
G
j
[…]
for further processing in GCN layers.
The subsequent procedure follows the same steps as in TP-GCN, including GCN processing, concatenation, and final classification.
5.5 Hyperparameter Configuration and Optimization Criteria
We propose four architectures that take advantage of the GCNConv and GraphConv models, resulting in eight distinct hypermodels. Each hypermodel features a set of hyperparameters that are automatically tuned based on the characteristics of the specific dataset.
Key hyperparameters for both convolutional models include the number of GCN and fully connected layers, unit configurations, activation functions, and optional mechanisms like batch normalization and dropout. The architectures also permit modifications in pooling operations, L1 regularization, learning rate schedules, optimizer selection with associated parameters, batch size, and loss functions. For the GraphConv model, we optimize neighborhood aggregation methods due to the absence of an adjacency matrix. In the two TE-GCN hypermodels, embedding dimensions are adjusted. The specific ranges and types of tuned hyperparameters are detailed in Table 1. For all models, we meticulously handled masked values (represented by 1-1- 1) to mitigate biases introduced by irrelevant entries in specific columns of the node vectors.
Table 1: Hyperparameters and Their Tuning Ranges/Types
Hyperparameter Range
Graph Convolutional Layers
Number of layers 1-5
Hidden Units 16-512
Skip Connection Y/N
Dropout flag: Y/N; rates: 0.2-0.7
Batch norm flag: Y/N; momentum: 0.1-0.999; eps:1e-5-1e-2
Activation ReLU, Leaky_ReLU, ELU, Tanh, Softplus, GELU
GraphConv Aggr add, mean, max
Pooling Method mean, add, max
Fully Connected Layers
Number of layers 1-3
Dense Units 16-512
Dropout flag: Y/N; rates: 0.2-0.7
Batch norm flag: Y/N; momentum: 0.1-0.999; eps:1e-5-1e-2
Activation ReLU, Leaky_ReLU, ELU, Tanh, Softplus, GELU
Optimizer and Learning Rate Scheduler
• Optimizer
Learning Rate 1e-5-1e-2 (log)
Weight Decay 0-1e-3
L1 0-1e-3
Type of Optimizers
Adam $\beta_{1}$: 0.85-0.99; $\beta_{2}$:0.99-0.999
SGD momentum: 0.0-0.9
RMSprop $\alpha$: 0.9-0.999; momentum: 0.0-0.9; eps: 1e-9-1e-7
• Learning Rate Schedulers
Step step size: 1-50; $\gamma$: 0.1-0.9
Exponential $\gamma$: 0.85-0.99
Reduce-on-Plateau factor: 0.1-0.9; patience: 1-50; threshold: 1e-4-1e-2; eps: 1e-8-1e-4
Polynomial power: 0.1-2; total_iters: 2-300
Cosine Annealing T_max: 10-100; eta_min: 1e-6-1e-2
Cyclic base: 1e-5-1e-2 (log); max: 1e-3-1e-1 (log); step_size_up: 5-200
One Cycle max: 1e-3-1e-1; total_steps: batch_size*1000; pct_start: 0.1-0.5
Loss Function CrossEntropy, MultiMargin
Batch Size 16, 32, 64, 128, 512
Embedding Dim 10-50
  • 1.
    Note: The abbreviations for parameters and hyperparameters in this table are derived from the PyTorch library, such as “ep” for epsilon [52], ensuring clarity and consistency in terminology.
In selecting the optimal model and hyperparameter configuration, the criteria varied on the basis of data set type (balanced or imbalanced). For balanced datasets, models were ranked according to test accuracy, with ties broken by the standard deviation of test loss to ensure stability; if both metrics were equal, the model with the lowest test loss was selected. This approach maximizes performance across all classes while promoting consistent generalization. For imbalanced data sets, the models were ranked by the weighted F1 score, which balances precision and recall and is appropriate to evaluate the performance of minority classes, often underrepresented by precision metrics. In the event of tied F1 scores, the test loss was considered, followed by the loss standard deviation for additional ties. Early stopping and pruning were used to prevent overfitting and improve computational efficiency.
6 Experiment
6.1 Data Description and Preprocessing
To evaluate model performance under different data conditions, we selected two dataset types: highly imbalanced and well-structured balanced datasets. Specifically, we used three different benchmarks: the Patients data set for the imbalanced condition and the BPI12-A and BPI12-O [60] data sets for the balanced conditions.
The Patients data set is a synthetic data set that comprises 2,142 case (sequence) IDs, each representing individual patient interactions and processes within the healthcare system. It contains a variety of activities (events) along with attributes at both the node (event) and graph (sequence) levels. In particular, this data set exhibits a significant class imbalance in six outcome categories: the majority class makes up 40.74% of the cases, while the minority class constitutes just 1.12%, giving a ratio of roughly 36:1. At the graph level, the dataset includes three numeric attributes and one categorical attribute, while the node-level data includes one universal categorical attribute, three numerical attributes, and one specific categorical attribute for select nodes. This dataset was chosen to assess GCN models due to its complex attribute structure.
The BPIC12-A and BPIC12-O datasets contain traces of the application processes for personal loans and overdrafts, respectively, within a multi-national financial institution. Each data set includes one numeric attribute at the graph level and two universal categorical attributes across all nodes. Both data sets were curated to achieve a balanced distribution between outcomes, with BPIC12-A containing 2,224 cases per outcome and BPIC12-O 802 cases per outcome. Each trace is classified into one of three outcomes —“approved (accepted)”,“declined” or “canceled” - corresponding to the final activity in the trace. These balanced data sets support robust evaluation of model performance across outcome categories.
Our model pipeline accommodates both imbalanced datasets (where traces share a final activity but vary in outcomes) and balanced datasets (with distinct final activities). By leveraging node and graph level attributes, this approach enhances predictive accuracy across different outcomes and adapts effectively to different data structures.
All data sets contain recorded start and completion times for each event. We used start times to calculate edge weights. In the Patients dataset, a duration attribute —calculated as the difference between start and completion times— is included in node attributes. To compute our pseudo-embedding, event durations were rounded to the nearest minute, with durations under 5 minutes assigned individual bins, and those above 5 minutes distributed across 24 quantile-based uniform bins. In contrast, BPIC12-A/O datasets have zero durations across nodes, so no duration attribute was incorporated in the node attributes nor in the pseudo-embedding matrix. For all datasets, the activity attribute $\mathcal{A}_{i}$ was decomposed into two categorical variables (verb and description) as proposed in [59], enhancing the representation of node-level data by breaking key attributes into distinct components. This decomposition enables for more granular feature extraction, improving the model’s capacity to detect intricate data patterns.
6.2 Experiment Setup
For the Patients dataset, all hypermodel types (GCNConv and GraphConv) were applied. TP-GCN was excluded from the BPIC12A/O datasets due to the absence of a pseudo-embedding matrix. The GCN hypermodels were tuned with the Optuna optimization algorithm [61], set to maximize performance over 200 trials, each trial consisting of up to 300 epochs with a patience level of 30. An 80/20 train/validation split was used for each class. Following the tuning phase, optimal hyperparameters were extracted from the best trial, allowing direct retrieval of the best-performing model with these parameters. Training continued for the full 300 epochs to assess improvements beyond the identified best epoch.
7 Results
In this section, we evaluate the performance of four distinct architectures based on GCNConv (G) and GraphConv (C) models, applied across both imbalanced and balanced datasets: One-Level Input (O-G/C), Two-Level Input (T-G/C), Two-Level Embedding (TE-G/C), and Two-Level Pseudo-Embedding (TP-G/C).
7.1 Models with Imbalanced Dataset
Table 2: Classification Report for Each Class of GCN Models
C O-GCNConv O-GraphConv T-GCNConv T-GraphConv TP-GCNConv TP-GraphConv TE-GCNConv TE-GraphConv S
0 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 92
1 0.7838 1 0.8788 0.7953 0.9828 0.8792 0.7838 1 0.8788 0.7838 1 0.8788 0.7793 0.9943 0.8737 0.7768 1 0.8744 0.7838 1 0.8788 0.7838 1 0.8788 174
2 1 1 1 0.625 1 0.7692 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 1 5
3 0.9545 1 0.9767 1 0.9048 0.95 1 0.9048 0.95 0.9524 0.9524 0.9524 0.9091 0.9524 0.9302 1 1 1 1 1 1 1 0.9524 0.9756 21
4 0.8649 1 0.9275 0.8611 0.9688 0.9118 0.9143 1 0.9552 0.8889 1 0.9412 0.9091 0.9375 0.9231 0.9333 0.875 0.9032 0.8649 1 0.9275 0.9143 1 0.9552 32
5 1 0.4808 0.6494 0.931 0.5192 0.6667 0.9636 0.5096 0.6667 0.9808 0.4904 0.6538 0.9815 0.5096 0.6709 0.9643 0.5192 0.675 1 0.4904 0.6581 0.9815 0.5096 0.6709 104
A 0.8738 0.8692 0.8762 0.8738 0.8715 0.8738 0.8762 0.8785 428
M 0.9339 0.9135 0.9054 0.8687 0.8959 0.8628 0.9436 0.9024 0.9084 0.9343 0.9071 0.9044 0.9298 0.899 0.8997 0.9457 0.899 0.9088 0.9414 0.9151 0.9107 0.9466 0.9103 0.9134 428
W 0.8998 0.8738 0.859 0.8853 0.8692 0.8581 0.8969 0.8762 0.8639 0.8968 0.8738 0.8599 0.8945 0.8715 0.8595 0.8956 0.8738 0.8627 0.902 0.8762 0.8622 0.9012 0.8785 0.8662 428
  • $\bullet$
    C: Class; S:Support; A: Accuracy; M: Macro Average F1; W: Weighted Average F1;
  • $\bullet$
    For each model, columns are precision, recall and F1-score, respectively.
Table 3: Hyperparameter Matrix for GCNConv and GraphConv Models
Model F1 B G(L) G(U) G(A) SC G(BM) G(BE) G(D) G(M) P D(L) D(U) D(A) D(BM) D(BE) D(D) Opt LR WD Sch Loss L1
O-G(I) 0.859 32 3 224 tanh F max 1 200 ELU 0.6507 3.839e-3 0.2456 RMS 1.857e-6 6.954e-5 Step MM 4.066e-5
(0.0103) (142) 102 tanh F 0.4504 2.13e-3 (0.9080, (28,
(0.8738) 187 GELU T 0.2662 4.145e-3 0.6204, 2.888e-8) 0.4188)
O-C(I) 0.8581 32 3 96 l_rl F 0.7746 7.298e-4 max mean 1 193 ELU 0.4741 5.322e-4 RMS 6.028e-8 6.985e-4 Exp MM 2.013e-5
(0.0376) (77) 170 ReLU T max (0.9464, (0.8956)
(0.8692) 123 l_rl F 0.5253 4.217e-4 mean 0.6985, 9.437e-8)
T-G(I) 0.8639 32 2 88 GELU F max 2(S) 69 ReLU 0.1601 Adam 1.248e-3 7.736e-4 Cos CE 3.424e-4
(0.0395) (88) 151 l_rl T 0.1764 2.471e-3 133 GELU 0.9650 9.671e-3 0.3444 (0.8837, (3.037e-3,
(0.8762) 1(C) 188 GELU 0.2754 8.581e-3 0.3062 0.9405) 50)
T-C(I) 0.8599 32 4 148 l_rl T 0.9083 6.797e-3 mean max 1(S) 53 GELU 0.2254 RMS 8.200e-6 9.062e-4 RP CE 2.017e-5
(0.0616) (42) 82 GELU T 0.8783 6.551e-3 add 3(C) 131 sp 0.4915 6.877e-3 (0.9478, (Max, 0.2471,
(0.8738) 250 ELU T 0.6116 3.451e-3 0.2922 add 85 tanh 0.8631, 4, 9.555e-3,
54 ELU T 0.6215 2.972e-3 add 215 ReLU 0.1363 2.086e-3 3.593e-8 ) 8.474e-6)
TP-G(I) 0.8595 32 2(N) 245 GELU add 1(S) 228 l_rl Adam 1.361e-4 8.761e-4 OCL CE 7.324e-5
(0.0526) (61) 203 GELU 0.5817 9.520e-3 2(C) 91 l_rl 0.1195 (0.9374) (1.831e-2,
(0.8715) 4(P) 38 GELU 0.5377 2.934e-3 0.1011 232 ELU 0.4174 9.586e-3 0.1100 0.9305) 0.1203,
194 ReLU 54000)
196 ELU
159 sp 0.2075 6.602e-3 0.4605
1(C) 32 ReLU T
TP-C(I) 0.8627 64 2(N) 194 tanh mean add 2(S) 243 ReLU 0.1235 Adam 7.155e-4 8.389e-4 Exp CE 1.577e-4
(0.0564) (86) 152 GELU 0.2008 3.538e-3 add 52 ReLU 0.5230 7.547e-3 (0.8965, (0.9602)
(0.8738) 1(P) 119 GELU add 2(C) 112 ReLU 0.6955 2.701e-3 0.9399)
3(C) 164 GELU F 0.7264 5.268e-3 max 219 sp 0.2987
120 l_rl F 0.2337 1.258e-4 max
TE-G(I) 0.8622 128 5(E) 162 l_rl 0.4643 2.321e-3 add 1(S) 156 GELU 0.2540 RMS 1.425e-5 1.567e-4 Poly MM 2.207e-7
(0.0260) (100) (11) 72 l_rl 1(C) 181 tanh (0.9008, (119,
(0.8762) 222 ELU 0.6657 6.918e-3 0.2334, 1.1527)
97 l_rl 0.7483 4.479e-3 0.1012 4.253e-8)
161 ReLU 0.3695 4.693e-3
5(N) 147 tanh 0.4960
241 tanh 0.6208 2.883e-3 0.3867
57 l_rl 0.0287 6.263e-4
162 tanh
184 l_rl
1(C) 213 l_rl T
TE-C(I) 0.8662 64 1(E) 199 GELU 0.8902 4.456e-3 0.1767 max mean 1(S) 117 sp Adam 2.589e-3 5.502e-4 Cos CE 1.791e-4
(0.0467) (98) (12) 222 l_rl mean 2(C) 119 sp (0.9420, (83,
(0.8785) 2(N) 128 l_rl 0.2398 max 89 l_rl 0.9709) 2.605e-3)
1(C) 130 l_rl T 0.6489 9.365e-3 max
  • $\bullet$
    Model: G: GCNConv; C: GraphConv
  • $\bullet$
    A: Accuracy; F1: F1(loss std)(Acc); B: Batch size (Best Epoch); G(L)/D(L): Number of hidden G(GCN)/D(Dense) layers; G(U)/D(U): Units ;G(A)/D(A): Activation functions; G(BE)/D(BE): Batch normalization epsilon; G(BM)/D(BM): Batch normalization momentum; G(D)/D(D): Dropout rates; SC: Skip Connection flag: G(M): Aggregation method for GraphConv; P: Pooling Method; Opt:Optimizer; LR: Learning Rate; WD: Weight Decay; Sch: Learning Rate Scheduler; Loss: Loss function; L1: L1 regularize; Empty cell in (BM)/(BE)/(D): No batch normalization or dropout applied.
  • $\bullet$
    *:(N): Node input layer (E): Embedding layer; (P):Pseudo-embedding input layer; (S): Graph level input layer; (C): Concatenation GCN/Dense layer; $\dagger$: Embedding dimensions
  • $\bullet$
    l_rl: Leary_ReLU; sp:softplus; CE: CrossEntropy; MM: MultiMargin.
  • $\bullet$
[…]
Figure 1: Learning Curves of GCN Models Applied on Imbalanced Dataset
7.1.1 Overall Performance
Table 3 shows the classification reports for each model on the highly imbalanced data set, detailing precision, recall, and F1 score metrics for individual classes. The optimized hyperparameters are summarized in table 4, along with accuracy and loss standard deviation during training. The learning curves for each model, retrained over 300 epochs, are shown in Figure 1, providing information on convergence behavior and overfitting. In this Section we perform a comprehensive analysis of these results, comparing different model performances and discussing the underlying factors contributing to the observed performance differences.
Across all models, accuracy remains relatively consistent, with only minor variations in F1 scores, largely attributable to the dataset’s imbalanced nature. Interestingly, two-level input models (separating node and graph features) generally outperform one-level input models, where features are aggregated at the node level. Furthermore, GCNConv models demonstrate superior stability and F1 scores with simpler two-level inputs, whereas GraphConv models excel with more complex inputs, such as those incorporating embeddings.
The stability of GCNConv models stands out, demonstrated by consistently lower test loss standard deviations and smoother learning curves compared to GraphConv models. Notably, the O-GCNConv models show the most consistent performance (lowest loss std 0.0103), likely due to their simpler structure. Adding embedding inputs of key node features to GCNConv models not only avoids instability, but also reduces loss variance, suggesting enhanced robustness. The TE-GraphConv model, in particular, achieves higher F1 scores and a lower standard deviation, showing the benefits of embedding-based inputs in GraphConv architectures. This trend is mirrored between the T-GraphConv and TP-GraphConv models, where pseudo-embedding inputs boost both stability and performance. In general, TE and TP variations underscore the importance of embedding inputs to optimize GraphConv models, especially with unbalanced datasets.
The learning curve plots confirm that all models were well fine-tuned, demonstrating high initial F1 scores and low loss across epochs. However, fluctuations in certain models’ learning curves highlight the challenges of the imbalanced dataset, particularly in learning from the minority class. The minimal gap between training and validation loss, where validation loss is often lower, suggests good generalization to unseen data despite dataset imbalances. During hyperparameter optimization, we employed early stopping to identify the optimal epoch, allowing us to determine the point of peak performance. In some instances, overfitting occurred, indicating a decline in generalization ability with further training. Upon retraining the best model with optimal hyperparameters, we observed a slight decrease in the F1 score, typically within 1% to 3%, consistent with expected variability due to the stochastic nature of the training process.
7.1.2 Class-Specific Performance
Across all models, recall scores for classes 0 through 4 consistently exceed 0.875, with 75% of these classes achieving a perfect recall of 1. This reflects the effective classification of the majority classes on the part of all models. In contrast, class 5 exhibits an unusually low recall of approximately 0.5, with notable variability among models; for example, O-GraphConv and T-GraphConv achieve precision and recall values of 0.4808 and 0.5096, respectively. These inconsistencies indicate potential problems with the quality or distribution of the data for this class. First, the presence of noisy, mislabeled, or irrelevant features could hinder the models’ ability to generalize effectively. If the training samples for class 5 contain inconsistencies, the models may struggle to establish appropriate decision boundaries. Another issue arises if the feature representations for class 5 are overly similar to those of other classes. High similarity may impede the models’ ability to differentiate class 5 effectively. Finally, if the distribution of test data for class 5 deviates from that of the training data, the models may under-perform during evaluation, resulting in lower recall and precision metrics.
Among all classes, class 0 achieves perfect precision, recall, and F1 scores (all 1s), likely due to its higher sample count (92), which facilitates accurate generalization across models. In contrast, class 1 performance shows minor variations between models, with precision values clustering around 0.78, while most models demonstrate high recall rates (greater than 98%), with the majority scoring 1. This indicates that the models are effective in accurately identifying genuine instances of class 1. Furthermore, an examination of the confusion matrix reveals that, across the eight models, 44 to 48 instances of class 5 are consistently misclassified as class 1, contributing to one of the lowest precision values for this class. This observation underscores the data quality issues inherent in class 5, which appear to affect all models uniformly.
In addition, class 2 presents a notable challenge due to its limited sample size of only five instances, hindering the models from achieving balanced precision and recall. While all other models attain perfect scores (all 1s), the O-GraphConv model exhibits lower performance, with a recall of 0.625 and a precision of 0.7692, indicating the model’s limitations in effectively handling underrepresented classes.
Despite their relatively smaller sizes compared to classes 0 and 1, classes 3 and 4 demonstrate satisfactory performance overall, with many models achieving perfect recall scores. However, we note that 3 to 5 instances of class 5 are misclassified as class 4, resulting in a decrease in precision. In light of these data quality issues, it is important to recognize that in domains where minority classes are critical, such as healthcare, a trade-off between performance criteria may be necessary during fine-tuning. Specifically, enhancing generalization capabilities for minority classes may lead to a decline in accuracy for majority classes. This intentional imbalance can represent a strategic compromise in applications where the accurate detection of minority classes is paramount.
7.1.3 Hyperparameters
Our analysis of hyperparameter settings highlights the critical role of precise tuning in achieving robust performance in graph neural networks. Using Optuna’s Bayesian optimization, we efficiently explored the hyperparameter space to identify optimal configurations for each model, thus mitigating the computational burden and manual effort typically associated with traditional grid search methods.
The choice of optimizer and learning rate is crucial in fine-tuning performance, with Adam and RMSprop yielding the best results. In GraphConv models, lower learning rates facilitated smoother convergence and reduced oscillations, while higher rates in GCNConv models accelerated learning in simpler architectures without sacrificing stability, affirming their robustness. Furthermore, the combination of smaller batch sizes, ranging from 32 to 128, with carefully tuned learning rates and weight decay rates significantly enhances model performance in imbalanced datasets. Smaller batch sizes enable more frequent weight updates, which is particularly beneficial for capturing nuanced patterns within minority classes. This strategy, when paired with optimized learning rates from a relatively broad spectrum ($6\times 10^{-8}$ to $2\times 10^{-3}$), provides a robust hyperparameter searching process that strikes a balance between rapid convergence and stability, facilitating effective learning from underrepresented samples. Furthermore, the relative lower weight decay values, ranging from $6\times 10^{-5}$ to $9\times 10^{-4}$, serve as a regularization mechanism that confirms the success of the tuning procedure in preventing overfitting while preserving critical features associated with minority classes. Overall, the synergy among smaller batch sizes, adaptive learning rates, and appropriately tuned weight decay underscores the effectiveness of our hypermodels in achieving this beneficial parameter combination, enhancing gradient propagation and exploration of the loss landscape for improved generalization across all classes in imbalanced datasets.
Interestingly, our findings indicate that simply increasing the number of hidden layers and units does not consistently correlate with improved model performance. For example, the O-GCNConv model, which utilizes three hidden layers and 224 units, surpasses the performance of the TE-GCNConv model, which comprises five hidden layers and 162 units. This phenomenon suggests the existence of an optimal complexity threshold for graph neural networks, where additional layers or units contribute diminishing returns. This observation is consistent with the principle of Occam’s Razor in model design, indicating the importance of achieving a balance between model complexity and predictive efficacy.
Skip connections (SCs) provide substantial advantages in managing imbalanced data within our GCN models, with architectures such as T-GCNConv and TE-GCNConv consistently outperforming those without SCs. This effectiveness can be attributed to three key factors, all of which align with our successful hypermodel tuning design. First, SCs enhance gradient flow, stabilizing training, and mitigating challenges such as vanishing or exploding gradients, which is particularly beneficial when learning from sparse minority class samples. Second, SCs help preserve essential features across layers, ensuring that minority class characteristics are maintained, thus enhancing model generalization. Third, SCs facilitate balanced feature learning, reducing overfitting on majority classes and promoting robust optimization through shorter gradient paths. These factors collectively reinforce the efficacy of our hypermodel tuning, demonstrating how thoughtful architecture choices can significantly improve model performance in imbalanced datasets.
7.1.4 Recommendations for Model Selection
For imbalanced data, it is essential to select models that perform robustly on both minority and majority classes. Two-level input models consistently outperform one-level models in handling class imbalance, with GCNConv-based models showing stronger generalization and stability, as reflected in their superior F1 scores and recall metrics across datasets. For conservative classification, GraphConv-based models with embedding inputs and early stopping provide comparable recall with added precision, notably in TE-GraphConv, which helps minimize false positives. The pseudo-embedding approach enhances performance by incorporating temporal relations, though it is sensitive to matrix variations. For cases involving embeddings or complex input structures, GraphConv with early stopping may be advantageous because of its stronger performance than GCNConv in such settings. TE-G(C) and TP-G(C) models consistently achieve higher recall, making them suitable for tasks where missing minority class predictions is costly. In contrast, T-G(C) models prioritize precision, making them ideal when false positives have significant consequences.
7.2 Models with Balanced Datasets
Table 4: Hyperparameter Matrix of GCNConv and GraphConv Models for BPI(A/O) Datasets
Model A B G(L) G(U) G(A) SC G(BM) G(BE) G(D) G(M) P D(L) D(U) D(A) D(BM) D(BE) D(D) Opt LR WD Sch Loss L1
O-G(A) 1 32 2 188 ELU T 0.4679 6.673e-3 0.3680 max 3 171 ReLU 0.4045 1.212e-3 0.2913 Adam 4.236e-2 6.323e-3 Exp MM 4.557e-4
(18) 85 ELU T 0.6117 5.621e-3 0.4486 109 ELU 0.1688 3.723e-3 0.4262 (0.8863, (0.9441)
182 ELU 0.9120)
O-G(O) 1 32 2 158 tanh T 0.0281 2.547e-4 0.4896 add 2 130 l_rl 0.4479 9.357e-3 0.3127 SGD 1.964e-3 3.927e-3 Cy MM 6.112e-4
(13) 216 ELU T 0.9216 5.337e-3 0.3833 180 l_rl 0.3666 9.1e-4 (0.0868) (7.214e-3, 9.382e-2, 36)
O-C(A) 1 32 3 211 ReLU F 0.1597 add mean 1 189 ReLU 0.6328 7.977e-3 0.1636 Adam 9.233e-4 6.863e-3 Poly MM 9.753e-5
(21) 100 sp F 0.0507 6.398e-4 0.1852 add (0.8887, (129,
75 ELU F 0.4934 add 0.9167) 0.5612)
O-C(O) 1 32 5 111 tanh F 0.5339 9.711e-3 0.3955 max max 1 210 sp 0.3473 2.428e-3 0.3955 Adam 2.223e-3 3.593e-3 Step MM 7.391e-5
(20) 128 sp T add (0.9081, (20
124 ReLU F add 0.9283) 0.7002)
44 GELU F 0.3403 max
71 l_rl T mean
T-G(A) 1 64 1 169 GELU T 0.6317 2.408e-3 max 2(S) 41 l_rl RMS 4.547e-3 8.561e-3 Cos MM 5.908e-4
(29) 105 l_rl (0.9154, (51,
1(C) 34 l_rl 0.8371, 9.053e-8) 6.309e-4)
T-G(O) 1 512 3 154 sp F 0.6040 5.235e-3 0.195 max 3(S) 156 sp 0.6054 8.760e-3 0.407 Adam 3.951e-4 7.458e-3 RP MM 9.109e-4
(16) 238 ReLU T 0.438 53 ReLU 0.7485 2.799e-3 0.221 (0.975, (Max, 0.5128,
163 l_rl F 0.0144 9.534e-3 111 ELU 0.3305 2.911e-3 0.926) 16, 4.604e-4
1(C) 120 ReLU 0.3416 2.631e-3 0.316 7.449e-5)
T-C(A) 1 32 2 151 tanh F 0.2745 1.149e-3 0.469 mean max 1(S) 111 tanh RMS 7.234e-4 4.415e-3 Cy MM 5.292e-4
(25) 128 sp F 0.4512 6.046e-3 max 1(C) 169 l_rl 0.243 (0.9111, 0.7603, 8.523e-8) (5.0392, 25, 0.0019)
T-C(O) 1 64 5 197 ReLU F 0.2854 add add 3(S) 247 ReLU 0.0678 2.667e-3 Adam 9.804e-3 5.032e-3 OCL MM 1.449e-5
(25) 98 ELU F 0.3419 5.304e-3 add 124 l_rl 0.8186 6.601e-3 0.1938 (0.8851, (0.0879,
162 l_rl F max 0.9008) 0.449)
172 ReLU F max 1(C) 227 ELU
191 GELU F 0.3995 3.511e-3 0.4485 max
TE-G(A) 1 64 2(E) 48 l_rl add 2(S) 48 GELU 0.3663 Adam 2.393e-3 2.099e-3 RP CE 3.930e-4
(31) (16) 237 ELU 0.6045 8.165e-3 0.2205 128 GELU (0.9692, (Max,
5(N) 128 ReLU 0.1164 7.970e-3 1(C) 83 ReLU 0.3612 0.9394) 0.8104,
192 GELU 0.5109 8.158e-3 23,
228 l_rl 0.1853 8.930e-3 7.53e-3,
201 ReLU 1.620e-5)
116 ELU
1(C) 128 GELU T 0.2656
TE-G(O) 1 64 1(E) 125 GELU 0.2835 8.908e-3 0.2889 max 3(S) 103 sp 0.6693 3.278e-3 RMS 1.907e-4 8.362e-3 Poly MM 8.788e-4
(17) (13) 50 ReLU 93 GELU (0.9021, (93,
4(N) 180 ReLU 0.0103 2.119e-3 95 ELU 0.7441, 1.1485)
164 GELU 0.2460 1(C) 114 ELU 0.4481 9.331e-8)
66 l_rl 0.3423 6.425e-3 0.1233
4(C) 145 GELU T 0.7255 8.292e-3 0.1577
114 tanh T
202 ReLU T 0.6147 5.549e-3 0.2826
73 l_rl T 0.2547 6.275e-3 0.2660
TE-C(A) 1 128 4(E) 235 ELU add mean 3(S) 131 ReLU 0.1011 Adam 4.579e-4 2.018e-3 RP MM 3.315e-4
(23) (18) 111 l_rl 0.2249 max 215 GELU (0.9477, (Max,
230 l_rl mean 194 tanh 0.3789 8.481e-3 0.9033) 0.8993,
105 GELU 0.5153 6.239e-3 0.1027 max 1(C) 156 ReLU 0/5428 9.075e-3 1,
4(N) 123 GELU 0.3155 6.208e-3 0.2360 mean 1.042e-4,
62 l_rl max 9.602e-5)
92 GELU 0.2555 mean
119 GELU 0.7320 7.155e-3 mean
1(C) 59 tanh T 0.3805 4.642e-3 max
TE-C(O) 1 32 5(E) 34 sp 0.3921 6.222e-3 mean mean 2(S) 130 ELU SGD 1.165e-3 7.960e-3 OCL MM 8.182e-4
(31) (21) 192 l_rl 0.6985 2.935e-3 0.2289 max 156 ELU (0.6997) (29.12e-2,
48 l_rl add 2(C) 204 ReLU 61,000,
111 l_rl add 246 sp 0.2550 7.048e-3 0.3087 0.2517)
232 l_rl 0.0729 9.9516e-4 0.1193 mean
4(N) 176 GELU 0.3532 9.875e-3 0.2951 mean
98 ELU mean
243 sp add
213 ELU 0.0218 9.528e-3 max
1(C) 207 sp F add
  • $\bullet$
    The table inherits the abbreviation from Table. 4
7.2.1 Overall Performance
The hyperparameter matrix in Table 4 and the learning curves in Figure 2 provide a comprehensive view of the performance of the model in the well-balanced BPI12 A/O datasets. We omit the confusion matrix, as all models achieve perfect precision, recall, and F1 scores (i.e., 1.0) across all classes, rendering the matrix redundant. All models demonstrate well-tuned performance, and the GCNConv models exhibit greater stability than the GraphConv models, consistent with previous findings on unbalanced datasets. The rapid convergence of certain models suggests that they can quickly adapt to the well-balanced datasets, achieving optimal performance with relatively few iterations (between 13-31 epochs). Although all models reach perfect accuracy given the characteristics of the data set, the focus of this study extends beyond the maximization of accuracy. In particular, certain models also achieve exceptional performance in minority classes on unbalanced datasets, highlighting the success of our fine-tuning procedure. This setup provides a valuable opportunity to benchmark how different hyperparameter strategies influence model efficiency, training dynamics, and generalization. The uniform accuracy across models allows us to isolate hyperparameter effects, offering deeper insights into the factors driving model behavior.
Figure 2: Learning Curve Of GCN HyperModels on Balanced Datasets
7.2.2 Hyperparameters
GCN Layers
A primary distinction between GCNConv and GraphConv models lies in their architecture depth and node-level input handling. First, GraphConv models often rely on deeper structures to capture intricate node feature relationships, though this can reduce stability. Conversely, GCNConv models, such as T-GCNConv and O-GCNConv, achieve similar performance with shallower architectures, underscoring their efficiency. Second, GCNConv models process embedding inputs more effectively, requiring fewer layers to capture essential node attributes, while GraphConv models need added depth to achieve comparable results. Interestingly, when processing non-key node attributes, both models benefit from deeper layers, as these features lack strong discriminative power. Skip connections are notably prevalent in GCNConv models, helping in training stability and gradient flow, especially in deeper networks. This feature allows GCNConv models to maintain performance without requiring excessive depth, resulting in enhanced stability and robustness over GraphConv models. In summary, GCNConv models demonstrate superior stability and efficiency with simpler architectures, while GraphConv models require added complexity to reach similar performance.
In examining activation functions across GCN layers, we find that Leaky_ReLU and GELU are preferred over softplus and $tanh$, suggesting that their effective gradient flow and non-linearity were optimized by our tuning, enhancing performance in both balanced and imbalanced datasets. This consistency across dataset types underscores our hypermodels’ robustness in supporting stable learning dynamics regardless of class distribution. Embedding models notably apply lower dropout rates in GCN layers, which appears to support generalization while minimizing overfitting. Additionally, minor variations in epsilon and momentum across models suggest these parameters have limited performance impact during tuning, while well-calibrated dropout rates in embedding models further reinforce stability.
Dense Layer
Our analysis reveals that both single-level (O) and two-level (T) input models have shallower dense layer structures following GraphConv layers compared to GCNConv layers. For O models, this suggests that GraphConv requires fewer dense layers for effective feature capture, probably because of its efficient feature propagation. In T models, this shallowness reflects the integration of graph-level attributes, reducing the need for deeper dense layers. This finding underscores that GraphConv models can leverage graph-level information more effectively, reducing the complexity of feature extraction compared to GCNConv. In particular, the depth of dense layers processing concatenated features from both input levels is unified at one, emphasizing GraphConv’s efficiency in handling graph-level information in balanced datasets. However, in embedding-based models, the depth difference in dense layers is less pronounced, indicating that the structure of the dense layer is influenced more by input complexity than by the embedding strategy. These results highlight the flexibility of embedding-based models in the dense layer configuration, potentially due to the additional feature abstraction offered by embeddings.
Another interesting finding regards the choice of activation functions in dense layers, where ReLU is more frequently employed than Leaky_ReLU and GELU, especially compared to GCN layers. This preference for ReLU likely reflects its simplicity and efficiency in adding non-linearity, which can enhance the model’s ability to capture complex patterns without excessive computational overhead. Moreover, this choice aligns with the relatively straightforward structure of graph-level input data, suggesting an effective tuning approach that balances complexity with model performance.
In particular, the BPI12O dataset requires deeper GCN layers with higher dropout rates and denser layer structures, reflecting its smaller size and higher variability in node attributes. This necessitates stronger regularization to maintain stability during training, indicating that the tuning procedure effectively addresses the dataset’s complexity.
Optimizer and Learning Rate
The Adam optimizer is more frequently selected over RMSprop for balanced data sets in models, probably due to its adaptive learning rate, which enhances performance in simpler and more stable data structures. Adam’s dynamic adjustment of learning rates aligns well with balanced datasets, supporting smoother and more precise optimization. In terms of learning rate schedulers, no specific trend emerged between datasets or models, underscoring the need to tailor the schedule to the unique characteristics of each dataset and model. This variety reflects a robust tuning strategy that optimizes model convergence by matching the optimizer and scheduler to the needs of the dataset. For example, Adam, when combined with schedulers such as Exponential Decay or Cosine Annealing [62], strikes an effective balance between exploration and fine-tuning. In contrast, models that use SGD with Cyclical learning rates [63] leverage greater exploration, enhancing generalization to more complex or noisy data distributions.
7.2.3 Recommendations for Model Selection
Given the perfect overall performance of all models on balanced data, and considering that all models are fine-tuned with early stopping and regularization, model selection should focus on the structure of inputs. We recommend choosing models that efficiently utilize multilevel inputs, such as two-level or pseudoembedding-based models, to leverage complex relationships and enhance generalization capabilities. If stability is a concern, the simpler architecture of T-GCNConv makes it a better option.
8 Conclusion and Discussion
This study presents HGCN(O), a toolkit for predicting event sequence outcomes in PBPM using self-tuned GCN models. Our results show that graphs can effectively represent temporal sequence data, including the full and partial overlap of concurrent activities, and achieve superior prediction performance. They also show that learning the hyperparameters of graph representations and processing layers is worthwhile, as it can facilitate several downstream tasks, such as node classification and outcome prediction. Our approach encodes different elements of event sequences into graph representations and develops hyper-GCN models with dynamically tuned hyperparameters, suitable for both highly unbalanced and well-balanced datasets. In turn, our graph representation maps graph entities to low-dimensional vectors while preserving the graph structure and inter-node relationships. Specifically, our toolkit brings together four different GCN architectures, using two GCN layer types (GCNConv and GraphConv) and different input structures. O-GCN models seamlessly integrate event- and sequence-level attributes at the node level into a single input. T-GCN models separate event (node) and sequence (graph) attributes into distinct inputs. TP-GCN models build on T-GCN by incorporating pseudo-embeddings. TE-GCN models extend T-GCN with embeddings of key event attributes. The architecture, hyperparameters and evaluation metrics of each model are automatically optimised to ensure performance and stability under different conditions. Our results show that T-GCN models with GCNConv layers excel at handling unbalanced datasets, while all models achieve high accuracy on balanced datasets, highlighting the importance of flexible model selection based on input structure and stability requirements. In addition, embedding structures in GraphConv models shows modest performance gains when coupled with early stopping, again highlighting the benefits of adaptive techniques for improved generalisation. Future studies can build on our findings by exploring alternative GCN architectures, input structures, and hyperparameter configurations. In particular, we plan to further investigate the impact of different shallow embedding techniques, such as Node2Vec and DeepWalk, on GCN performance. Finally, we note that shallow models learn their embeddings by representing graph entities as vector points in a latent Euclidean space. However, graphs encoding sequential process data may have irregular structures and different shapes, so the Euclidean space may not be adequate to represent the graph structure [7]. We plan to extend our approach to handle node embeddings that could be defined as a continuous density, using autoencoders to learn the appropriate divergence with respect to the Gaussian distribution.